# Sijia Liu2,3, Mingyi Hong1

### v1

@@ B. Hemagglutinin and nucleoprotein
Table 8: Examples of generated text from different unlearning methods on the MUSE-News dataset, comparing outputs under two decoding configurations $(T,p)=(0.2,0.2)$ and $(0.2,1.0)$. Each case shows the worst response among $128$ generations. Failed unlearning is indicated by undesired answers in red, successful unlearning by green.
@@
Table 9: Examples of generated text from different unlearning methods on the MUSE-News dataset, comparing the worst outputs under decoding configuration $(T,p)=(0.8,1.0)$ using $k=1$ and $128$ generations.
For MUSE-Books, the unlearning task is inherently less challenging, as all methods achieve zero knowledge memory on the forget set under deterministic decoding (see Table 3 in reisizadeh2025blur), leading to negligible information leakage across settings. As shown in Table 10, we report $\widehat{\text{leak@$k$}}$–RS under both low- and high-randomness decoding configurations, $(T,p)=(0.2,0.2)$ and $(T,p)=(1.0,1.0)$. Consistent with expectations, GradDiff demonstrates complete robustness, while SimNPO and BLUR-NPO reveal only marginal traces of residual knowledge leakage.
Table 10: Evaluation of $\widehat{\text{leak@$k$}}$–RS on MUSE-Books across different numbers of generations $k$ under low- and high-randomness settings, $(T,p)=(0.2,0.2)$ and $(1.0,1.0)$. Values are shown for $(T,p)=(0.2,0.2)$ $\mid$ $(T,p)=(1.0,1.0)$, corresponding to the low- and high-randomness settings, respectively.
Method $k=1$ $k=2$ $k=4$ $k=8$ $k=16$ $k=32$ $k=64$ $k=128$
SimNPO $0\mid0$.04 0.00 $\mid$ 0.07 0.00 $\mid$ 0.08 0.00 $\mid$ 0.11 0.00 $\mid$ 0.12 0.00 $\mid$ 0.14 0.00 $\mid$ 0.15 $0\mid0$.16
BLUR-NPO $0\mid0$.01 0.00 $\mid$ 0.01 0.00 $\mid$ 0.02 0.00 $\mid$ 0.04 0.00 $\mid$ 0.05 0.00 $\mid$ 0.07 0.00 $\mid$ 0.09 $0\mid0$.10
NPO-SAM $\mid0.00$ 0.00 $\mid$ 0.00 0.00 $\mid$ 0.00 0.00 $\mid$ 0.00 0.00 $\mid$ 0.01 0.00 $\mid$ 0.01 0.00 $\mid$ 0.03 $0\mid0$.05
NPO $\mid0.00$ 0.00 $\mid$ 0.00 0.00 $\mid$ 0.00 0.00 $\mid$ 0.01 0.00 $\mid$ 0.01 0.00 $\mid$ 0.01 0.00 $\mid$ 0.02 $0\mid0$.03
GradDiff $\mid0.00$ 0.00 $\mid$ 0.00 0.00 $\mid$ 0.00 0.00 $\mid$ 0.00 0.00 $\mid$ 0.00 0.00 $\mid$ 0.00 0.00 $\mid$ 0.00 $\mid0.00$
WMDP. We report $\widehat{\text{leak@$k$}}$–Acc for RMU and NPO on WMDP-bio. For this core metric, we check whether the chosen option for each generation of the unlearned model matches the WMDP ground truth.
Figure 5: $\widehat{\text{leak@$k$}}$–RS for RMU model on the WMDP dataset using Zephyr-7B-beta model. Rows correspond to different pairs of $(T,p)$.
Fig. 5 reports the $\widehat{\text{leak@$k$}}$–Acc of the RMU model across multiple numbers of generations $k$ and sampling parameters $(T,p)$ pairs. Similar to the TOFU and MUSE-News benchmarks, the parameter $p$ has a stronger influence than $t$ on information leakage (see the rows corresponding to (T,p)=(1.0,0.2)(T,p){=}(1.0,0.2) and (T,p)=(0.2,1.0)(T,p){=}(0.2,1.0)). However, for the NPO model, $\widehat{\text{leak@$k$}}$–Acc remains flat across all $k$ and $(T,p)$ pairs, matching the greedy decoding score, $26\%$fan2024simplicity. The flat $\widehat{\text{leak@$k$}}$–Acc for NPO arises from overly aggressive forgetting, which disrupts answer selection.
We further validate our observation for $\widehat{\text{leak@$k$}}$–Acc using LLM-as-Judge (LJ) metric. Since WMDP is originally designed for answer selection, we extend its evaluation to free-form generation by comparing model outputs with the description of the correct option. This generation-based evaluation provides a more informative assessment than accuracy, as it enables analysis of the model’s reasoning traces and thus better captures potential information leakage in LLMs. Details of the prompt design are provided in Appendix B.1.
Fig. 6 illustrates that RMU exhibits a clear information-leakage trend under high-randomness sampling $(T,p)=(1.0,1.0)$ and nearly no information leakage under low randomness $(T,p)=(0.2,0.2)$, consistent with the pattern observed for $\widehat{\text{leak@$k$}}$–Acc. The NPO method generates gibberish responses (see Table 11), resulting in $\widehat{\text{leak@$k$}}$–LJ scores near 0, indicating no observable information leakage. Consistent with our findings using $\widehat{\text{leak@$k$}}$–Acc, the NPO method remains robust under $\widehat{\text{leak@$k$}}$–LJ evaluation.
As our results show, for the TOFU and MUSE-News benchmarks across all models and for RMU on the WMDP dataset, $\widehat{\text{leak@$k$}}$$\sf{CoreM}(\cdot,\cdot)$] increases sharply with $k$ using various core metrics $\sf{CoreM}(\cdot,\cdot)$, indicating that generating more samples under relatively high probabilistic decoding sharply raise the chance of leakage. Moreover, top-$p$ emerges as the primary parameter governing this leakage behavior.
@@ D. United States
Table 11: Example responses using NPO model on WMDP dataset. The first example shows that the model replaces sensitive information with “/******/”, while in the second example it degenerates into incoherent text, just gibberish. These behaviors demonstrate that NPO effectively removes target knowledge, serving as a robust unlearning method on WMDP.
Our findings reveal a key weakness of current unlearning methods, they remain vulnerable to decoding strategy and highlights the need for more robust approaches. Here, we propose and implement a simple extension of the NPO algorithm, denoted as NPO-Fix, which augments the forget set with detected leakage instances identified through generation-based evaluation. While built on the same token-level NPO framework, NPO-Fix introduces a dynamic dataset augmentation step that integrates generation into the unlearning process, aiming to mitigate sequence-level leakage. As shown in Table A2, NPO-Fix improves over the vanilla NPO method on the TOFU dataset; however, leakage still grows with $k$, confirming that this simple fix remains insufficient and that more principled solutions are needed. Further implementation details are provided in Appendix F.
4 Conclusion
We showed that existing unlearning methods appearing successful under greedy decoding evaluations, continue to leak sensitive information under realistic probabilistic decoding. To quantify this leakage, we introduced leak@$k$, a semantic and distributional meta-metric that captures worst-case responses across multiple generations. Our systematic evaluation on TOFU, MUSE, and WMDP demonstrates that current unlearning methods consistently leak across a wide range of temperature and top-$p$ settings. These results highlight the urgent need for new approaches that achieve reliable forgetting while preserving overall model utility.
Appendix
Appendix A Related Work
LLM Unlearning. Due to the amount of training data of LLMs, retraining LLMs from scratch is infeasible. Hence, it is critical to exploit LLM unlearning techniques. LLM unlearning is typically formulated as a regularized optimization problem, where a penalty term on the retain loss is added to the forget objective. The challenges of choosing proper losses, especially forget loss imply new complexities in capturing the optimal balance between unlearning and utility. To address this, several approaches have been proposed, including gradient ascent (GA)[thudi2022unrolling, yao2023large, maini2024tofu], NPO [zhang2024negative], and SimNPO [fan2024simplicity]. Recently, [bu2024unlearning] and [reisizadeh2025blur] studied LLM unlearning through the lens of multi-task optimization and simple bi-level optimization, respectively.
Evaluating Unlearning. Evaluating unlearned models requires metrics that capture whether they avoid reproducing sensitive information from the forget set while still generating accurate and useful responses for the retain set. Various metrics from natural language generation have been adapted for LLM unlearning, including ROUGE-L [lin-2004-rouge], BERTScore [zhang2019bertscore], cosine similarity [cer-etal-2017-semeval], and entailment-based scores [ferrandez2008te4av, yao2024accurate, poliak-2020-survey]. ROUGE-L measures lexical overlap between the generated response and the ground truth. BERTScore computes cosine similarity between contextual embeddings of the generated and reference texts, using pre-trained BERT representations to capture semantic alignment and robustness to paraphrasing. Cosine similarity applied directly to sentence embeddings (e.g., from models like Sentence-BERT [reimers-2019-sentence-bert]) provides a lightweight semantic measure, though it is less fine-grained than token-level BERTScore. Finally, entailment scores from natural language inference (NLI) models assess whether the generated output entails or contradicts the reference, thus evaluating factual and logical consistency.
Probabilistic Evaluation. Current LLM unlearning methods rely on greedy decoding for evaluation, thereby overlooking LLMs’ probabilistic nature [maini2024tofu, shi2024muse, li2024wmdp]. Recent works advance probabilistic evaluation to more accurately assess the robustness of unlearning methods. [scholten2024probabilistic] demonstrate that greedy decoding fails to capture the full output distribution of LLMs and thus introduces probabilistic analysis metrics for a more comprehensive evaluation. In their work, they employ entropy optimization and adaptive temperature scaling to incorporate distributional information into the unlearning process. Similarly,  [yuan2024closer] point out the unreliability of greedy decoding by introducing token entropy as an evaluation metric and proposes maximizing prediction entropy on the forget set to make the generation distribution approximate random guessing. While our metric leak@$k$ is also based on probabilistic evaluation, our contributions go further in three key directions. (1) Semantically, current probabilistic analysis operates at the statistical level, without grounding in semantics or task-level meaning, whereas leak@$k$ explicitly incorporates semantic evaluation. (2) Methodologically, prior approaches rely on single-shot decoding, with entropy capturing only pointwise uncertainty; in contrast, leak@$k$ evaluates at the distributional level, assessing whether the entire output space continues to contain forgotten knowledge. (3) Analytically, we conduct comprehensive experiments showing that leak@$k$ yields consistent results across diverse unlearning benchmarks using both generation-based and answer-selection evaluations, whereas the entropy-based methods they propose fail to provide such reliability (see Section 3).
Appendix B Evaluation Details
B.1 LLM Judge
Prompt Construction. To ensure the reliability of the LLM Judge as an evaluator, we design a standardized system prompt that explicitly defines the evaluation task. Specifically, the LLM is provided with the model-generated response, the gold-standard answer, and the original query, and is required to output a clear verdict, indicating whether the candidate response successfully captures the essential information contained in the reference answer. Following prior unlearning evaluation [wang2025reasoning], we adopt a three-category classification prompt for WMDP dataset to enhance the robustness of judgment. This framework enables the LLM Judge to more effectively identify both explicit information leakage and implicit knowledge retention. The full prompt template used for this evaluation is presented below.
{Question + Correct Answer + Model Generation}
Assistant:
{Score: 1–3 only}
{Question + Gold Answer + Model Generation}
Assistant:
{Score: YES or NO}
Judge Selection To ensure accurate and efficient detection of potential information leakage, selecting an appropriate LLM as the judge is crucial. Among the available models, the GPT series by OpenAI [achiam2023gpt] has been widely adopted as evaluation judges due to their strong language comprehension and reasoning abilities.
For TOFU benchmark, we employ GPT-4o-mini  [hurst2024gpt] as the judge model. Since TOFU primarily consists of self-narrative tasks that are relatively simple, GPT-4o-mini provides sufficient reasoning capacity to handle them effectively, achieving both high accuracy and efficiency.
In contrast, for the WMDP dataset, we adopt the GPT-o3-mini model [openai_gpto3_mini_2025]. This WMDP dataset involves complex biomedical knowledge, which requires the judge to possess extensive pretraining knowledge and deeper reasoning capabilities. GPT-o3-mini offers these strengths, making it a more suitable evaluator for challenging biomedical reasoning tasks.
Below, we present examples from WMDP and TOFU dataset to illustrate our evaluation process.
B.2 Entailment Score
To quantitatively assess semantic consistency between the generated response and the gold answer, we employ an entailment-based evaluation metric inspired by natural language inference (NLI). Specifically, we use a pretrained NLI model implemented via the transformers library [wolf2019huggingface], utilizing the sileod/deberta-v3-base-tasksource-nli  [^1] https://huggingface.co/sileod/deberta-v3-base-tasksource-nlicheckpoint. This model determines whether the generated response semantically entails the gold answer. Given a pair of texts—the model-generated response (as the premise) and the gold answer (as the hypothesis)—the classifier outputs a binary decision: an entailment label of 1 indicates that the generated response contains the essential information of the gold answer (i.e., information leakage), while 0 denotes the absence of entailment. This approach provides a systematic and reproducible way to detect whether the model’s output preserves knowledge that should have been unlearned.
We apply the ES on the TOFU benchmark to evaluate semantic-level information leakage. The TOFU dataset consists of factual question–answer pairs centered on autobiographical information about specific authors. Leveraging the NLI model’s capability for deep semantic understanding, ES determines whether a model’s generated response implicitly or explicitly entails the gold factual answer. This evaluation approach extends beyond surface lexical overlap and instead captures semantic consistency, thereby providing a more robust indication of whether sensitive factual knowledge has been retained. The following example illustrates how ES is applied in practice on the TOFU benchmark.
Appendix C State-of-the-Art Unlearning Methods
LLM unlearning deals with two objectives: the forget loss, that aims to remove the influence of undesirable information from the model, and the retain loss, which ensures that the model’s overall utility is preserved. The retain loss is typically formulated using cross-entropy (CE), or alternatively with the RMU loss developed in [li2024wmdp], given by
$\bullet\ \ell_{\text{GA}}$ [maini2024tofu, thudi2022unrolling] treats the forget set as negative examples and directly maximizes their log-likelihood, driving the model’s predictions to diverge from them.
$\bullet\ \ell_{f}\!=\!\ell_{\text{NPO},\beta}$ for a given $\beta\!\geq\!0$ [zhang2024negative], which penalizes the model when it assigns a higher likelihood to forget examples relative to a reference model $\bm{\theta}_{0}$.
$\bullet\ \ell_{f}\!=\!\ell_{\text{SimNPO},\beta,\alpha}$ for given $\beta,\alpha\!\geq\!0$ [fan2024simplicity] removes the dependence on a reference model and normalizes by sequence length, introducing a reward margin $\alpha$ to adjust forgetting strength.
$\bullet\ \ell_{f}=\ell_{\text{RMU},f}$ [li2024wmdp] perturbs hidden representations, pushing them toward a random direction $\mathbf{u}$ so that information from the forget set cannot be reliably recovered.
The corresponding losses are given as follows:
LLM unlearning problems are typically formulated as a regularized optimization problem [liu2022continual, yao2023large, maini2024tofu, eldan2023whos, zhang2024negative] (which leverage some weighted some of forget and retain objectives) or some forms of bi/multi-objective optimization problem [reisizadeh2025blur] [bu2024unlearning] (which enforces some kind of priorities among the loss functions). Within the regularized formulation, various algorithms correspond to specific choices of retain and forget loss pairs, GradDiff uses ((A1), (A3)), NPO uses (N/A, (A4)), SimNPO uses ((A1), (A5)), and RMU uses ((A2), (A6)). Also, BLUR–NPO is a proposed method based on the bi-level formulation [reisizadeh2025blur] using the retain loss in (A1) and the forget loss in (A4).
C.1 Entropy Optimization Unlearning
In our TOFU evaluation, we include a probabilistic, NPO+ENT. Here, we provide the technical details of NPO+ENT, the entropy-regularized unlearning method proposed in [scholten2024probabilistic]. This method aims to control the uncertainty of the model’s output distribution during the unlearning stage by introducing an additional entropy loss. This loss minimizes the entropy of the token distribution, encouraging the model to produce less diverse outputs and concentrate probability mass, thereby reducing the likelihood of generating undesired responses. Formally, the NPO+ENT objective is defined as
where $\ell_{\text{CE}}(y\mid x;\bm{\theta})$ and $\ell_{\text{NPO},\beta}(y\mid x;\bm{\theta})$ are provided in (A1) and (A4), respectively. Here, $\lambda$ is a weighting coefficient balancing the contribution of the entropy term. The entropy loss for a given pair $(x,y)$ is given by
$\ell_{\text{ENT}}(y\mid x;\bm{\theta})=\frac{1}{m}\sum_{t=1}^{m}H\!\left(\pi(y_{t}\mid y_{<t},x;\bm{\theta})\right),$
where $\pi_{\theta}(y_{t}\mid y_{<t},x)$ denotes the predictive distribution over the vocabulary at time step $t$, $m$ is the sequence length, and $H(q)=-\sum_{i=1}^{|V|}q_{i}\log q_{i}$ is the entropy function. In our experiment for TOFU daataset, we set $\lambda\!=\!1$ and the parameter $\beta\!=\!0.5$ for $10$ epochs with a learning rate of $1\!\times\!10^{-6}$.
Appendix D Unbiasedness of $\widehat{p}_{k}(\tau)$
Fix a threshold $\tau\in[0,1]$ and let $p_{\tau}:=\mathop{\rm Pr}\nolimits(S\geq\tau)$. For the $n$ i.i.d. draws, define indicators $Y_{i}:=\mathbf{1}\{s_{i}\geq\tau\}$, so $Y_{1},\ldots,Y_{n}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathrm{Bernoulli}(p_{\tau})$, and $c_{\tau}=\sum_{i=1}^{n}Y_{i}$ counts the number of “successes”. Let $I=\{I_{1},\ldots,I_{k}\}$ be a uniformly random $k$-subset of $\{1,\ldots,n\}$ (independently of $Y$). Conditional on the realization of $Y$, the probability that all $k$ chosen elements are failures equals the hypergeometric term $\mathop{\rm Pr}\nolimits\big(Y_{I_{1}}=\cdots=Y_{I_{k}}=0\,\big|\,Y\big)=\frac{\binom{n-c_{\tau}}{k}}{\binom{n}{k}}$. Taking expectation over $Y$ (law of total expectation), we get
$\mathbb{E}\!\left[\frac{\binom{n-c_{\tau}}{k}}{\binom{n}{k}}\right]=\mathop{\rm Pr}\nolimits\big(Y_{I_{1}}=\cdots=Y_{I_{k}}=0\big).$
By exchangeability of the i.i.d. indicators, the joint distribution of $(Y_{I_{1}},\ldots,Y_{I_{k}})$ is the same as that of $(Y_{1},\ldots,Y_{k})$ (now viewed as an ordered $k$-tuple of distinct indices). Hence, we have
since the $Y_{i}$ are independent with $\mathop{\rm Pr}\nolimits(Y_{i}=1)=p_{\tau}$. Therefore, we can write
$\mathbb{E}\!\left[1-\frac{\binom{n-c_{\tau}}{k}}{\binom{n}{k}}\right]=1-(1-p_{\tau})^{k}=p_{k}(\tau),$
so $\widehat{p}_{k}(\tau)$ is an unbiased estimator of $p_{k}(\tau)$. Finally, by linearity of expectation, we get
Appendix E Additional Evaluation Results and Details
E.1 Additional Evaluation Results
As shown in Fig. A1, even at a relatively high value of p=0.8p{=}0.8 with a low T=0.2T{=}0.2, no considerable information leakage is observed for TOFU benchmark across all models.
Fig. A2 illustrates the performance of unlearned models on the retain set for TOFU dataet across different $(T,p)$ configurations. As observed, when either $T$ or $p$ remains low, model utility is well preserved, and the performance drop from low- to high-randomness decoding is negligible.
A B C D E
Original 23.2% 19.6% 23.3% 21.5% 12.4%
NPO 1.3% 0.9% 1.3% 2.4% 94.2%
RMU 9.6% 8.6% 10.0% 9.3% 62.5%
Appendix F A Simple Attempt to Mitigate Information Leakage
We have shown that existing SOTA unlearning methods exhibit substantial information leakage with our metric, leak@$k$. The SFT-based unlearning methods (e.g., NPO) operate at the token-level, optimizing cross-entropy against reference labels. However, true knowledge leakage emerges at the sequence-level, where sensitive information can be reconstructed across multiple tokens. Capturing this requires semantic evaluation of full generations, beyond token-level probabilities. Therefore, we propose a simple dynamic training approach that integrates generation during unlearning. In the following, we outline the key steps of our proposed algorithm, which we denote as NPO-Fix.
(1) Baseline Unlearning. We train the target model $\bm{\theta}(0)$ with NPO loss provided in (A4) on the forget set $\mathcal{D}_{f}$ for $t_{0}$ iterations to obtain an initial unlearned model $\bm{\theta}(t_{0})$.
(2) Leakage Detection. For each $x\in\mathcal{D}_{f}$, sample candidate outputs $y\sim\pi(\cdot\mid x;\bm{\theta}{(t_{0})})$ under a probabilistic decoding. If $\text{CoreM}(y,y_{f})\geq\tau$ for the reference $y_{f}$, record $(x,y)$ as a leakage instance.
(3) Forget Set Expansion. Collect all leakage instances into an augmented forget dataset, given by $\tilde{\mathcal{D}}_{f}=\mathcal{D}_{f}\cup\{(x,y)\mid\text{CoreM}(y,y_{f})\geq\tau\}$.
(4) Iterative Refinement. Re-train the (unlearned) model $\bm{\theta}(t_{0})$ using the NPO loss on $\tilde{\mathcal{D}}_{f}$ for $t_{1}$ iterations to get the final unlearned model $\bm{\theta}(t_{0}\!+\!t_{1})$.
We conduct an experiment on the TOFU dataset. In Step (1), we train for $t_{0}=10$ epochs. In Step (2), each question in the forget set is prompted $32$ times using probabilistic decoding with temperature $T=1.0$ and top-$p=1.0$. We adopt ES as the core evaluation metric and set $\tau=1$ to construct the augmented dataset $\tilde{\mathcal{D}}_{f}$. In Step (4), the model obtained after Step (1) is further re-trained for $5$ epochs. Across all training phases, we use a learning rate of $1\times 10^{-6}$ and set the NPO hyperparameter $\beta=0.5$.
We evaluate NPO-Fix using $\widehat{\text{leak@$k$}}$–ES. As shown in Table A2, NPO-Fix achieves stronger unlearning performance than NPO and often even surpasses the Retrain baseline. However, $\widehat{\text{leak@$k$}}$–ES for NPO-Fix still shows substantial growth and remains high. Therefore, our dynamic dataset augmentation approach is not a complete solution to information leakage, but it highlights a promising direction for future unlearning algorithms: incorporating generation into the unlearning stage. Overall, our results emphasize that robust unlearning in LLMs is a non-trivial challenge requiring more principled advances.
Table A2: Comparison of NPO-Fix with other unlearning methods using $\widehat{\text{leak@$k$}}$–ES.
Method Number of Generations $k$
1 2 4 8 16 32 64 128
Retrain 16.9% 23.6% 30.6% 37.6% 44.1% 50.1% 55.7% 61.0%
Original 28.7% 40.5% 52.4% 63.1% 72.1% 79.4% 84.9% 89.1%
NPO 20.4% 29.2% 38.4% 47.3% 55.4% 62.5% 68.7% 74.3%
NPO-Fix 11.9% 17.9% 24.9% 32.5% 39.9% 47.2% 54.6% 61.9%

### v2

@@ E.4 MUSE Results
Table A7: Examples of generated text from different unlearning methods on MUSE-News, comparing outputs under two decoding configurations $(T,p)=(0.2,0.2)$ and $(0.2,1.0)$. Each case shows the worst response among $128$ generations. Failed unlearning is indicated by undesired answers in red, successful unlearning by green.
@@
Table A8: Examples of generated text from different unlearning methods on the MUSE-News dataset, comparing the worst outputs under decoding configuration $(T,p)=(0.8,1.0)$ using $k=1$ and $128$ generations.
E.5 WMDP results
Table  A9 shows the probability of LLMs choices with options E, illustrating NPO method seems to deviates LLM from making choices within A/B/C/D.
A B C D E
Original 23.2% 19.6% 23.3% 21.5% 12.4%
NPO 1.3% 0.9% 1.3% 2.4% 94.2%
RMU 9.6% 8.6% 10.0% 9.3% 62.5%
In Table  A10, we provide two examples of the NPO method on the WMDP benchmark. The model successfully masks sensitive information, replacing it with placeholders $\mathbf{/******/}$; it also frequently produces gibberish outputs, indicating a degree of over-unlearning that degrades general generation quality.
@@ D. United States
Table A10: Example responses using NPO model on WMDP dataset. The first example shows that the model replaces sensitive information with “/******/”, while in the second example it degenerates into incoherent text, just gibberish. These behaviors demonstrate that NPO effectively removes target knowledge, serving as a robust unlearning method on WMDP.
Appendix F RULE: Algorithm Description and Implementation Details
All unlearning experiments are conducted using LLaMA-3.2-1B as the base model on TOFU benchmark. Unlearning is performed on a single NVIDIA A40 GPU. We instantiate on four representative unlearning methods: NPO, GradDiff, SimNPO, and RMU.
Hyperparameter Configuration. For the experiments, the learning rate $\eta$ is set to $10^{-5}$ across all methods. The weighting coefficient $\beta$ is fixed to $0.5$ for NPO and $4.5$ for SimNPO. We set the forget and retain weight $\lambda=1.0$ for GradDiff, NPO, SimNPO, and RMU. For the SimNPO unlearning scheme, the auxiliary coefficient $\alpha$ is fixed to $1.0$. Further, for RMU, the steering coefficient $c$ is set to $2.0$. All hyperparameter configurations used in our experiments are summarized in Table A11.
Method $\eta$ $\beta$ $c$ $\alpha$ $\lambda$
GradDiff $10^{-5}$ 1.0
NPO $10^{-5}$ 0.5 1.0
SimNPO $10^{-5}$ 4.5 1.0 1.0
RMU $10^{-5}$ 2.0 1.0
Table A11: Hyperparameters for various unlearning methods on the TOFU benchmark.
RULE Configuration.
All models undergo an initial unlearning phase on the original dataset for $5$ epochs. At each subsequent generation stage, we randomly select $50\%$ of the questions from the dataset for both the forget set and the retain set to accelerate training. For each selected question, $20$ responses are generated. The generated responses are evaluated using ES, and only those exhibiting successful entailment are retained. Unlearning is then performed for $5$ epochs at each stage using the same hyperparameter configuration described above. This generation-incorporated unlearning process is repeated for a total of three iterations. Finally, we report a breakdown of the time cost associated with each major component of the RULE framework in Table A12.
Operational Stage Time / Iter (min) Total (min)
Initial Unlearning ($1\times$) $10$ $10$
Response Generation ($3\times$) 2020 $60$
Entailment Evaluation ($3\times$) 55 $15$
Enhanced Unlearning ($3\times$) $10$ $30$
Overall Total $\sim$ $2h$
Table A12: Time cost breakdown of the RULE unlearning framework. Generations are repeated for three iterations.